OpenStudy (anonymous):

Write the standard equation for the circle that passes through the points: (1, 7) (8, 6) (7, -1) I've gotten as far as having 100=-8d-6e. At this point I'm ready to find the values for d, e, and f. Buuuut problem is that I don't know how to solve for any of them since there are 2 variables in the equation. My teacher said substitution was sometimes the way to find the answers, but I did & I was wrong. I thought that d=-5 and e=-10. When I substituted the numbers into my equation, I got 100. However, it is still wrong. Where do I go from here?

OpenStudy (anonymous):

I also know that my general equation is \[x^2 + y^2 + dx + ey + f = 0\] so that is how I know I have to have an f value. Which I think is zero but I'm honestly not sure.

OpenStudy (badhi):

now they have given you three points that passes through the circle, (1,7), $$1+49+d(1)+e(7)+f=0\implies d+7e+f=-50 \quad (1)$$ (8,6) $$65+36+8d+6y+f=0\implies 8d+6e+f=101\quad (2)$$ (7,-1) $$49+1+7d-e+f=0\implies 7d-e+f=-50\quad (3)$$ did you get these three equations?

OpenStudy (anonymous):

The only thing different was #2 because I got \[8^2\] is 64 not 65 which when adding 36 would make it 100 and not 101. Otherwise I'm with you so far!

OpenStudy (mertsj):

second equation: 8d+6e+f=-100

OpenStudy (anonymous):

Yes, that is what I have for the second equation

OpenStudy (badhi):

sorry, a typo you're right. it has to be $$8d+6e+f=-100$$

OpenStudy (mertsj):

Turns out d = -8, e=-6, f=0

OpenStudy (anonymous):

But I need to know how to find it. But thank you for helping, too! But here is what I got when I subtracted the second equation from the first.. \[50=-7d+e\] Then the third from the second.. \[-50=d+5e\] Then the later from the first.. \[100=-8d-6e\]

OpenStudy (badhi):

your second equation is wrong isn't it? it has to be, $$-50=d+7e$$ you have to add the equations in a way that one variable disappears. that is multiply first with 7 and substract it with second equation. $$350+50=-49d-d\\ 400=-50d$$ d=-8

OpenStudy (anonymous):

I'm a little confused. I multiply which equation by 7 or both?

OpenStudy (anonymous):

Oh I see now. Which equation do I substitute d=-8 into?

OpenStudy (mertsj):

Any equation that contains only d and e

OpenStudy (anonymous):

Okay, thanks y'all for your help!!!

OpenStudy (mertsj):

yw